# Can AI Dream of Unseen Galaxies? Conditional Diffusion Model for Galaxy Morphology Augmentation

### v1

@@ 3 Data
To achieve this, we used the normal-depth Sloan Digital Sky Survey samples of Galaxy Zoo 2 (hereafter GZ2) [^2] https://data.galaxyzoo.org debiased by the method described in Hart et al. (2016), which contains 239,695 galaxy images labeled by human volunteers using a hierarchical morphological classification system (e.g., Lintott et al., 2008; Willett et al., 2013; Hart et al., 2016; Walmsley et al., 2023b). In this section, we describe the Galaxy Zoo 2 dataset, outline the preprocessing steps used to prepare the data for training the diffusion model, and explain the structure employed for galaxy morphology conditions. Despite we here focus on morphological features in this work, our method can be easily extended to other astrophysical properties, such as redshift, stellar or halo mass, dust content and so on, which we give more discussion in Section 5.
Figure 2: Upset plot to visulize the morphology distribution in our training samples. The gray dots and connections represents co-occurrence of different morphological features in the dataset. Bars indicate the counts of individual features and their combinations. We simply show primary tag categories, excluding detailed tags such as bulge prominence, the number of spiral arms, etc. This highly imbalanced distributions emphasize the necessity of incorporating synthetic data for robust machine learning training.
3.1 Galaxy Zoo Project
@@
As a citizen science project, Galaxy Zoo 2 benefits from volunteer contributes but such treatment also introduces noise. We here propriety training set quality before data size to ensure better performance of our generative model.
@@
3.2 Early-type Dust Lane Galaxy Catalog
In addition to the GZ2 galaxy morphology catalog, we also incorporate a catalog of early-type galaxies (ETGs) with prominent dust lanes from Kaviraj et al. (2012) to demonstrate our generative model’s capability to synthesize rare objects by extrapolating from common visual features and by enhancing downstream ML training pipelines.
Although early-type galaxies and dust lanes are individually common, their combination—“dusty” early-type galaxies (hereafter, D-ETGs) are rare, serving as a tracer of minor merger events and offering valuable insights into galaxy evolution. To date, Kaviraj et al. (2012) presents one of the largest D-ETG catalogs, containing 352 nearby D-ETGs ($0.10.01<z<$ 0.1) identified from approximately 300,000 galaxies in the GZ2 dataset. This catalog was constructed by first selecting galaxies with at least one volunteer marking the presence of a dust lane, which produced 19,000 galaxy candidates, and subsequently having these candidates visually inspected by expert astronomers to confirm both the presence of a dust lane and the galaxy’s early-type morphology.
We construct our rare object detection task by using this 352 D-ETGs as positive samples and demonstrate that our generative model can effectively generate synthetic D-ETGs to augment the training set and improve the performance of ML models in identifying these rare objects in Section 4.3. We excluded those 352 D-ETGs, as well as galaxies labeled as combination of early-type features (e.g., completely round or in-between round) and dust lane feature, when training our diffusion models. This exclusion was made to prevent potential data leakage and to highlight the capability of generative models in performing generative extrapolation on rare morphological structures.
4 Result
In this study, we address the data scarcity challenges in large-scale galaxy image analysis by leveraging diffusion models for data augmentation. We begin by evaluating the quality of the generated images and then demonstrate the effectiveness of the synthetic images in enhancing the performance of machine learning models across various scientific applications. These include general galaxy morphology classification and a rare object detection task focused on early-type galaxies with prominent dust-lane features, thereby underscoring the practical value of our methodology in real-world research environments.
Figure 3: Comparison of galaxy images generated by our model under various morphology-related text prompts, compared with real galaxy images. For each column annotated by a capital letter in top-right corner, the leading first image is a real one from SDSS, while the following three images are synthesized by our model. The annotated capital letters represents prompts used to generate simulated galaxies as Table 1 shows.
@@ 4 Result
K in-between round galaxy, {dust lane:1.3}
L edge-on galaxy, {dust lane:1.3}, with rounded edge-on bulge
@@
[…]
4.1 Generated Image Quality Evaluation
@@
Although these astrophysical metrics are closely related to the physical properties of galaxies, they are not well suited for our study for two reasons: (1) they may not capture the diverse and intricate visual details—such as spiral arms, mergers, dust lanes, lenses, or arcs—present in galaxy images, and (2) they cannot be directly applied to the pseudo-colored images used in our analysis. We also note here that there lacks unified evaluation system for synthesized or simulated galaxy images (e.g., Smith et al., 2022; Hackstein et al., 2023; Vičánek Martínez et al., 2024), and the metrics we use here pose new direction for future work.
Our evaluation framework assessed three critical aspects of generated galaxy images: (1) realism (visual fidelity relative to real observations), (2) diversity (morphological variety across outputs), and (3) consistency (adherence to input conditions). Formally, let the real image dataset be $\mathbf{X}=\{\mathrm{X}_{i}\mid i=1,\dots,N^{\prime},\;\mathrm{X}_{i}\in\mathbb{R}^{H\times W\times 3}\}$ and the synthesized dataset be $\mathbf{Y}=\{\mathrm{Y}_{j}\mid j=1,\dots,N,\;\mathrm{Y}_{j}\in\mathbb{R}^{H\times W\times 3}\}$. For a visual condition $p$ (e.g., GZ2-annotated features; see Section 3), conditional realism $\mathcal{R}^{p}_{C}$ is quantified as below:
@@
To evaluate the conditional consistency $\mathcal{C}^{p}$, we use the commonly adopted visual question answering (VQA) approahces (e.g., Hu et al., 2023; Cho et al., 2023; Astolfi et al., 2024). In detail, we train a series binary classifier $\mathbf{Q}_{i}$ tailored to different visual features, such as, the presence of spiral arm, if this is a ring on galaxy, etc. The conditional consistency $\mathcal{C}^{p}$ is then defined as the average accuracy of the classifiers $\mathbf{Q}_{i}$ under condition $p$:
@@
where $Q_{j}$ represents the number of binary classifiers employed for image $\mathrm{Y}_{j}$ ($11$ in this work), $A_{i}$ is the ground truth to classifier $\mathbf{Q}_{i}$ on image $\mathrm{Y}_{j}$, and $\mathbb{I}(\cdot)$ is the indicator function. The global consistency $\mathcal{C}$ is the average of $\mathcal{C}^{p}$ across all conditions $p\in\mathbf{P}$:
@@
Above metrics were tracked throughout the training iterations to monitor progressive improvements in image quality. As shown in Figure 4, the realism of the generated images rapidly improve during the first 1,000 steps before continuing to grow steadily. Examples in Figure 3 illustrates that our synthesized images closely resemble real galaxy images by capturing a wide range of morphological features—including spiral arms, dust lanes, and signs of mergers—according to the given conditions. Furthermore, Figure 4 reinforces that the diversity of the generated images increases continuously, demonstrating the model’s gradual mastery of detailed galaxy morphology, which is also evident in Figure 3.
Figure 4: The quantified consistency, diversity and realism of generated galaxy images with training steps increasing, under average prompts. The lower panel intuitively demonstrates that under various prompts, the more training steps evolve, the more real generated galaxy images are.
Figure 5: Pareto fronts showing the trade-off relationship between realism and diversity evaluation indicators of our diffusion model. The dashed lines are Pareto fronts representing the frontiers of optimal solutions to the three evaluation metrics. Different colors represent different types of morphological prompts as top legend shows.
We also notice that there is Pareto front between the realism and diversity of the generated images, as shown in Figure 5. This indicates that while our model can generate images with high realism, it may sacrifice some diversity in the process. This trade-off is common in generative models, where achieving high realism often leads to less diverse outputs. However, we find that our model maintains a good balance between these two aspects, producing images that are both realistic and diverse.
Overall, our diffusion model successfully generates high-quality galaxy images that closely resemble real observations. The model exhibits strong realism, diversity, and consistency, as evidenced by the visual fidelity, morphological variety, and adherence to input conditions of the generated images. These results underscore the potential of our approach to enhance machine learning performance in various scientific applications, as detailed in the following sections.
4.2 Improving Galaxy Morphology Classification with Synthetic Data
@@
As we have demonstrated that our diffusion model can generate high-fidelity galaxy images, we now show how these images can bring tangible scientific benefits. We first focus on classical binary classification tasks involving key galaxy morphologies, including spirals, edge-on disks, bars, and bulges. These features are critical for understanding various astrophysical processes, such as the role of spiral arms in star formation (e.g., Foyle et al., 2010; Schinnerer et al., 2017; Sun et al., 2024a), the close relationship between presence frequency of disk galaxies with $\Lambda$CDM cosmoloy (e.g., Neeleman et al., 2020; Nelson et al., 2023; Yan et al., 2024; Kohandel et al., 2024), the influence of bars on gas inflow and central starbursts (e.g., Friedli et al., 1994; Fraser-McKelvie et al., 2019; Lin et al., 2020; Chen et al., 2023), and the prominence of bulges in deciphering galaxy merger histories (e.g., Hopkins et al., 2010; Fontanot et al., 2011; Nedkova et al., 2024).
@@
While these morphological features are much common compared to those rare-events such as strong lensing, tidal features, and dust lanes, they are still challenging to identify due to the lack of high-quality labeled datasets and the inherent class imbalance. As our generative model can produce a large number of realistic galaxy images with specific morphological features, we can use these synthetic images to enhance the training set for machine learning models.
@@
In this experiment, we split the GZ2 dataset into a training set (90%) and a test set (10%). Specifically, we further randomly select 10% of the training data to form a validation set. For each task, we fix the number of negative samples at 10,000 and sample positive samples proportionally to maintain the same class distributions as the GZ2 full dataset (containing 239,695 galaxies). For the four selected morphological features, they have different difficulties in classification due to their different positive sample fractions and contamination rates from volunteer annotations. We here neglect the possible contaminations in the training set, as the contamination rate should not dominate the final results since we take a relative conservative data selection criteria as described in Section 3.
The positive fractions of the four features, as annotated by volunteers, are 0.43, 0.09, 0.10, and 0.07 for spiral, disk, bar, and bulge features, respectively. We then use the GZ2 dataset to train a convolutional neural network (CNN) model with a ResNet-18 backbone He et al. (2016) for each of the four morphological features. The model is trained using a binary cross-entropy loss function and an Adam optimizer with a learning rate of $10^{-4}$. We also apply standard data augmentation such as random cropping, rotation, and flipping to improve the model’s generalization ability.
@@
To evaluate the performance of our model, we use three metrics: purity, completeness, and F1-score. Purity measures the proportion of correctly classified positive samples among all samples predicted as positive, while completeness measures the proportion of correctly classified positive samples among all actual positive samples. The F1-score is the harmonic mean of purity and completeness, providing a balanced measure of the model’s performance. We compare our model’s performance with two baselines: (1) a pure supervised learning model trained only on the GZ2 dataset and (2) a model trained with focal loss Lin et al. (2017) to address class imbalance.
@@
The results are presented in Table 2. Our model outperforms both baselines across all four classification tasks, achieving the highest values for purity, completeness, and F1-score. For instance, in the spiral feature classification task, our model attains a purity of 0.69, completeness of 0.70, and F1-score of 0.70—surpassing the baseline model (0.67, 0.45, 0.54) and the focal loss model (0.65, 0.36, 0.46). Similar improvements are observed in the other three tasks. These findings demonstrate that our generative model effectively enhances the performance of machine learning models in identifying key morphological features within galaxy images, particularly for common galaxy morphology classification tasks.
@@
The success of our model stems from the high quality and diversity of the generated images, which provide a rich source of synthetic data for training downstream ML models. Since image generation is a more complex task than binary classification, generative models are less prone to overfitting on systematic biases in the data. This enables them to improve the generalizability of downstream ML models. In astrophysics, data distributions are intrinsically imbalanced due to physical processes governing cosmic evolution. However, robust ML models require large training datasets to avoid becoming trapped in local minima. Our method bridges this critical gap by generating synthetic data to address such imbalances.
@@
4.3 Enhancing Early-type Dust Lane Galaxy Detection via Extrapolation on Rare Morphologies

### v2

@@ III Data
To achieve this, we used the normal-depth Sloan Digital Sky Survey samples of Galaxy Zoo 2 (hereafter GZ2) [^2] https://data.galaxyzoo.org debiased by the method described in Hart et al. (2016), which contains 239,695 galaxy images labeled by human volunteers using a hierarchical morphological classification system (e.g., Lintott et al., 2008; Willett et al., 2013; Hart et al., 2016; Walmsley et al., 2023b). In this section, we describe the Galaxy Zoo 2 dataset, outline the pre-processing steps used to prepare the data for training the diffusion model, and explain the structure employed for galaxy morphology conditions. Despite we here focus on morphological features in this work, our method can be easily extended to other astrophysical properties, such as redshift, stellar or halo mass, dust content and so on, which we give more discussion in Section V.
Figure 2: Upset plot to visulize the morphology distribution in training samples. The gray dots and connections represents co-occurrence of different morphological features in the dataset. Bars indicate the counts of individual features and their combinations. We simply show primary tag categories, excluding detailed tags such as bulge prominence, the number of spiral arms, etc. This highly imbalanced distributions emphasize the necessity of incorporating synthetic data for robust machine learning training.
III.1 Galaxy Zoo Project
@@
As a citizen science project, Galaxy Zoo 2 benefits from volunteer contributes but such treatment also introduces noise. We prioritize training set quality over data size to ensure better performance of our generative model.
@@
III.2 Early-type Dust Lane Galaxy Catalog
In addition to the GZ2 galaxy morphology catalog, we also incorporate a catalog of early-type galaxies (ETGs) with prominent dust lanes from Kaviraj et al. (2012) to demonstrate GalaxySD’s capability to synthesize rare objects by extrapolating from common visual features and by enhancing downstream ML training pipelines.
Although early-type galaxies and dust lanes are individually common, their combination—“dusty” early-type galaxies (hereafter, D-ETGs) are rare, serving as a tracer of minor merger events and offering valuable insights into galaxy evolution. To date, Kaviraj et al. (2012) presents one of the largest D-ETG catalogs, containing 352 nearby D-ETGs ($0.01<z<0.1$) identified from approximately 300,000 galaxies in the GZ2 dataset. This catalog was constructed by first selecting galaxies with at least one volunteer marking the presence of a dust lane, which produced 19,000 galaxy candidates, and subsequently having these candidates visually inspected by expert astronomers to confirm both the presence of a dust lane and the galaxy’s early-type morphology.
We construct our rare object detection task by using this 352 D-ETGs as positive samples and demonstrate that GalaxySD can effectively generate synthetic D-ETGs to augment the training set and improve the performance of ML models in identifying these rare objects in Section IV.3. We excluded those 352 D-ETGs, as well as galaxies labeled as combination of early-type features (e.g., completely round or in-between round) and dust lane feature, when training conditional diffusion model. This exclusion was made to prevent potential data leakage and to highlight the capability of GalaxySD in performing generative extrapolation on rare morphological structures.
IV Result
In this study, we address the challenges of data scarcity in large-scale galaxy image analysis by leveraging diffusion models for data augmentation. We begin by evaluating the quality of the generated images and then demonstrate the effectiveness of the synthetic images in enhancing the performance of machine learning models across various scientific applications. These include general galaxy morphology classification and a rare object detection task focused on early-type galaxies with prominent dust-lane features, thereby underscoring the practical value of our methodology in real-world research environments.
Figure 3: Comparison of galaxy images generated by GalaxySD under various morphology-related text prompts, compared with real galaxy images. For each column annotated by a capital letter in top-right corner, the leading first image is a real one from SDSS, while the following three images are synthetic. The annotated capital letters represents prompts used to generate simulated galaxies as Table 1 shows.
@@ IV Result
K in-between round galaxy, {dust lane:1.2}
L edge-on galaxy, {dust lane:1.2}, with rounded edge-on bulge
@@
Table 1: Morphological conditions employed in generating the diverse galaxy images in Figure 3. In K and L prompts, {dust lane:1.3} assigns a $1.3\times$ weight to the dust lane feature during inference through cross-attention mechanism, emphasizing its prominence. See Appendix B for the detailed hierarchical morphology classification system.
IV.1 Generated Image Quality Evaluation
@@
Although these astrophysical metrics are closely related to the physical properties of galaxies, they are not well suited for our study for two reasons: (1) they may not capture the diverse and intricate visual details—such as spiral arms, mergers, dust lanes, lenses, or arcs—present in galaxy images, and (2) they cannot be directly applied to the pseudo-colored images used in our analysis. We also note here that there lacks a unified evaluation system for synthesized or simulated galaxy images (e.g., Smith et al., 2022; Hackstein et al., 2023; Vičánek Martínez et al., 2024), and the metrics we use here pose a new direction for future work.
Our evaluation framework assessed three critical aspects of generated galaxy images: (1) realism (visual fidelity relative to real observations), (2) diversity (morphological variety across outputs), and (3) consistency (adherence to input conditions). Formally, let the real image dataset be $\mathbf{X}=\{\mathrm{X}_{i}\mid i=1,\dots,N^{\prime},\;\mathrm{X}_{i}\in\mathbb{R}^{H\times W\times 3}\}$ and the synthesized dataset be $\mathbf{Y}=\{\mathrm{Y}_{j}\mid j=1,\dots,N,\;\mathrm{Y}_{j}\in\mathbb{R}^{H\times W\times 3}\}$. For a visual condition $p$ (e.g., GZ2-annotated features; see Section III), conditional realism $\mathcal{R}^{p}_{C}$ is quantified as below:
@@
To evaluate conditional consistency $\mathcal{C}^{p}$, we use the commonly adopted visual question answering (VQA) approaches (e.g., Hu et al., 2023; Cho et al., 2023; Astolfi et al., 2024). In detail, we train a series of classifiers $\mathbf{Q}_{i}$ tailored to different morphological features of galaxy images (hereafter, classifiers for VQA).
These classifiers for VQA adopt a ResNet18 architecture and correspond to 11 key morphological questions derived from the Galaxy Zoo 2 (GZ2) question tree as Figure B3 shown, covering 11 distinct visual features, with the test set accuracies listed in parentheses – (1) smoothness (80.8%), (2) roundness (91.2%), (3) edge-on disk orientation (94.5%), (4) bulge shape (86.6%), (5) bar presence (85.0%), (6) spiral arm presence (89.8%), (7) spiral arm tightness (69.0%), (8) number of spiral arms (63.5%), (9) bulge prominence (73.7%), (10) presence of odd features (69.4%), and (11) type of odd features (e.g., ring, merger, dust lane; 61.4%). These accuracies reflect the performance of each classifier on a test set of real galaxies that were completely excluded from training, which is defined as the sum of true positives across all classes divided by the total number of samples.
Overall, these performance indices indicate that these classifiers for VQA achieve high accuracy for visually salient features (e.g., disk orientation, spiral arms) but relatively lower performance for more ambiguous or fine-grained categories (e.g., number or type of odd features), reflecting the intrinsic variability of these morphological traits. Also, the limited performance of classifiers for VQA maybe affect the evaluation. However, we have also measured average conditional consistency on real galaxy images as a benchmark, which is 0.895, shown in Figure 4, which could enhance the reliability of consistency and provide a more robust basis for interpretation.
This benchmark value in real images below 1 indicates that the classifiers are not perfect and may occasionally misclassify even real images. However, this does not diminish the utility of conditional consistency. The classifiers achieve accuracies ranging from approximately 60% to 90%, which ensures that consistency can provide quantitative measurements of feature correctness. By using the consistency of real images as a reference benchmark, we can meaningfully interpret and compare the consistency of generated images. The performance improvements observed in downstream classification tasks (see Section IV.2) further validate this result.
Each classifier $\mathbf{Q}_{i}$ is designed to answer a specific morphological question as described above. They share the same ResNet18 architecture, and each classifier is trained from scratch independently, using standard cross-entropy loss and the Adam optimizer Kingma & Ba (2015) with a learning rate of $1\times 10^{-3}$. Training proceeds until convergence with early stopping based on the validation loss. Each classifier is trained on a class-balanced subset of real galaxy images from GZ2, labeled according to the corresponding answers to its target question. For example, the classifier for identifying bulge shape is trained on a dataset comprising galaxies annotated as having a rounded bulge, a boxy bulge, or no bulge, with an equal number of samples per class. The dataset is randomly split into 80% for training, 10% for validation, and 10% for testing.
@@
where $Q_{j}$ represents the number of classifiers employed for image $\mathrm{Y}_{j}$ (1111 in this work), $A_{i}$ is the ground truth to classifier $\mathbf{Q}_{i}$ on image $\mathrm{Y}_{j}$, and $\mathbb{I}(\cdot)$ is the indicator function, which returns $1$ if the condition inside is true, and 0 otherwise; it serves as a switch reflecting whether the stated condition holds. The global consistency $\mathcal{C}$ is the average of $\mathcal{C}^{p}$ across all conditions $p\in\mathbf{P}$:
@@
Above metrics were tracked throughout the training iterations to monitor progressive improvements in image quality. As shown in Figure 4, the realism of the generated images rapidly improve during the first 1,000 steps before continuing to grow steadily. Examples in Figure 3 illustrates that synthetic images by GalaxySD closely resemble real galaxy images by capturing a wide range of morphological features—including spiral arms, dust lanes, and signs of mergers—according to the given conditions. Furthermore, Figure 4 reinforces that the diversity of the generated images increases continuously, demonstrating the model’s gradual mastery of detailed galaxy morphology, which is also evident in Figure 3.
Figure 4: The quantified consistency, diversity and realism of generated galaxy images with training steps increasing, under average prompts. The benchmark line in the step vs. consistency figure is $y=0.895$, which denotes the consistency of real galaxy images under average prompts. The lower panel intuitively demonstrates that under various prompts, the more training steps evolve, the more real generated galaxy images are.
Figure 5: Pareto fronts showing the trade-off relationship between realism and diversity evaluation indicators of our conditional diffusion model. The dashed lines are Pareto fronts representing the frontiers of optimal solutions to the three evaluation metrics. Different colors represent different types of morphological prompts as top legend shows.
We also notice that there is a Pareto front between the realism and diversity of the generated images, as shown in Figure 5. This indicates that while GalaxySD can generate images with high realism, it may sacrifice some diversity in the process. This trade-off is common in generative models, where achieving high realism often leads to less diverse outputs. However, we find that our model maintains a good balance between these two aspects, producing images that are both realistic and diverse.
Overall, our fine-tuned conditional diffusion model, GalaxySD, successfully generates high-quality galaxy images that closely resemble real observations. The model exhibits strong realism, diversity, and consistency, as evidenced by the visual fidelity, morphological variety, and adherence to input conditions of the generated images. These results underscore the potential of our approach to enhance machine learning performance in various scientific applications, as detailed in the following sections.
@@
IV.2 Improving Galaxy Morphology Classification with Synthetic Data
[…]
@@ IV.2 Improving Galaxy Morphology Classification with Synthetic Data
While these morphological features are much common compared to those rare-events such as strong lensing, tidal features, and dust lanes, they are still challenging to identify due to the lack of high-quality labeled datasets and the inherent class imbalance. As GalaxySD can produce a large number of realistic galaxy images with specific morphological features, we can use these synthetic images to enhance the training set for machine learning models.
@@
For each binary classifier for classical tasks, we fix the number of negative samples at 10,000 and sample positive samples proportionally to maintain the same class distributions as the GZ2 full dataset (containing 239,695 galaxies), with the detailed proportions listed in corresponding parentheses. Specifically, the training set for each classifier consists of a positive sample size of 4,300 for spiral (43%), 900 for disk (9%), 1,000 for bar (10%), and 700 for bulge (7%) features and 10,000 negative samples, respectively. These different positive sample ratios lead to varying training set difficulties across classifiers, with lower positive fractions (e.g., bulge) generally resulting in harder classification tasks. We here neglect the possible contaminations in the training set, as the contamination rate should not dominate the final results since we take strict and conservative data selection criteria as described in Section III.
Specifically, we augment the training set by generating a large number of synthetic galaxy images based on specific morphological prompts. The synthetic-augmented training set for the “spiral” feature consists of 4,300 real galaxies (43%) and 5,700 synthetic ones as augmented positive samples; for the “disk”, “bar”, and “bulge” features, we use 900, 1,000, and 700 real images respectively, and augment them with synthetic images to reach 10,000 positive samples per task. In each case, 10,000 negative samples are randomly selected from real galaxies without the target feature, resulting in class-balanced training sets. Due to the limited number of real galaxy images from observation, the effective training set sizes of the synthetic-augmented and real datasets differ, according to the above description.
We evaluated the effectiveness of these augmentation-trained classifiers for classical tasks by comparing our model with two baselines: (1) a pure supervised learning classifier trained only on the GZ2 dataset and (2) a focal loss-trained classifier Lin et al. (2017) to address the class imbalance. Both methods are trained using only real images, with the same training set distribution as described in the previous paragraph, but without any synthetic augmentation.
The idea of the focal loss method is to modify the loss function, so that to address the issue of class imbalance and to improve the sensitivity of the model to hard-to-classify samples. Compared with the conventional cross-entropy loss, focal loss introduces a modulation mechanism that down-weights the contribution of easily classified examples while increasing the relative importance of misclassified ones. This property makes it particularly effective for tasks with skewed class distributions. The detailed formulation and implementation of focal loss are provided in the Appendix D.
In these experiments, we split the prepared dataset for each task into a training set (90%) and a test set (10%) and further randomly select 10% of the training data as a validation set. For both the real training set and the synthetic-augmented training set, we apply basic data augmentation operations, including rotation, cropping, flipping, zoom-in, and zoom-out, to enhance the diversity of the training samples and improve the model’s robustness to variations in galaxy orientations and scales. We then use the dataset to train a convolutional neural network (CNN) model with a ResNet18 backbone He et al. (2016) as the binary classifier for each of the four morphological features in baseline, focal loss and our synthetic-augmented methods.
These classifiers for classical tasks are trained using a binary cross-entropy loss function except the focal loss case, and the Adam optimizer with a learning rate of $1\times 10^{-3}$. All hyperparameters are kept identical across the three methods to ensure a fair comparison. Rather than training for a fixed number of epochs or focusing on training loss, we apply the early stopping criterion based on the validation loss, terminating training once no further improvement is observed. Across the four synthetic-augmented classifiers for classical tasks, convergence was reached within roughly a dozen epochs, depending on the specific task and dataset. The classifier for spiral galaxies converged in 12 epochs, the classifier for disk-shaped galaxies in 13 epochs, the classifier for barred galaxies in 14 epochs, and the classifier for galaxies with bulges in 14 epochs.
We note that these binary classifiers for classical tasks differ from those used for consistency evaluation of our generation model, primarily in their training data. The former use synthetic images from our model for data augmentation and simulate the class distribution of the GZ2 dataset, thus are not class-balanced. In contrast, the latter are trained solely on real galaxy images with balanced positive and negative samples to reliably identify specific morphological features.
@@
To evaluate the performance of these classifiers for classical tasks trained on our augmented dataset, we use three metrics: purity, completeness, and F1-score. Purity measures the proportion of correctly classified positive samples among all samples predicted as positive, while completeness measures the proportion of correctly classified positive samples among all actual positive samples. The F1-score is the harmonic mean of purity and completeness, providing a balanced measure of the classifiers’ performance.
@@
The results are presented in Table 2. Our method outperforms both baselines across all four classification tasks, achieving the highest values for purity, completeness, and F1-score. For instance, in the spiral feature classification task, our method attains a purity of 0.69, completeness of 0.70, and F1-score of 0.70—surpassing the baseline model (0.67, 0.45, 0.54) and the focal loss model (0.65, 0.36, 0.46). Similar improvements are observed in the other three tasks. These findings demonstrate that our generative model effectively enhances the performance of machine learning models in identifying key morphological features within galaxy images, particularly for common galaxy morphology classification tasks.
@@
The success of our method stems from the high quality and diversity of the generated images, which provide a rich source of synthetic data for training downstream ML models. Since image generation is a more complex task than binary classification, generative models are less prone to overfitting on systematic biases in the data. This enables them to improve the generalizability of downstream ML models. In astrophysics, data distributions are intrinsically imbalanced due to physical processes governing cosmic evolution. However, robust ML models require large training datasets to generalize well on real-data. Our method bridges this critical gap by generating synthetic data to address such imbalances.
The enhanced performance of our approach in real-world data classification tasks establishes a critical link between image quality evaluation—leveraging the heuristic metrics of realism, consistency and diversity outlined in Section IV.1—and scientific utility for research. This connection not only enables a more holistic assessment of synthesized images but also highlights the practical value of GalaxySD for advancing studies focused on galaxy morphology.
@@
IV.3 Enhancing Early-type Dust Lane Galaxy Detection via Extrapolation on Rare Morphologies